Sherry maneuvered her way through the first 60 miles in 4 hours. Then she doubled her speed for the last part of the trip. If the total trip was 300 miles,how many hours did it take Sherry to complete the last part?

Found 2 solutions by mananth, dkala99:
Answer by mananth(16946)

Sherry maneuvered her way through the first 60 miles in 4 hours. Then she doubled her speed for the last part of the trip. If the total trip was 300 miles,how many hours did it take Sherry to complete the last part?
First she went at 60 mph for 60 miles
Then she went at 120 mph for 240 miles
The time taken for the last part = Distance / rate
= 240 / 120= 2 hours

Answer by dkala99(12)

Total distance of the trip = 300 mi
Distance of the first trip = 60 mi
Total number of trips = 300/60 = 5 trips to complete the total trip
Total trip 300 mi = 60 mi + 60 mi + 60 mi + 60 mi + 60 mi

Speed for the first 60 mi in 4 hr = Distance/Time = 60/4 = 15 mph
Sherry doubled her speed for the last part of the trip.
speed for the last part of the trip = 2 * 15 mph = 30 mph
Time for the last part of the trip = Distance/Speed = 60/30 = 2 hrs